# Data for Mathematical Copilots: Better Ways of Presenting Proofs for Machine Learning

### v1

@@ 2 An Overview of Data-Related Issues
For the other category, datasets around formal mathematics frequently suffer from other issues than those listed above; for instance, datasets for formal mathematics often come at a range of difficulties, from competition-level problems (Zheng et al., 2022) to undergraduate textbooks (Azerbayev et al., 2023), to some results that are at the mathematical state-of-the-art (Scholze, 2022; Bordg et al., 2022). Moreover, formalized mathematics evaluation can be easily carried out in a safe and automated manner since proofs can be checked by the ITP.
@@
  • Tool misalignment: Various automation tools exist that aid in proving small, intermediate lemmas arising in the formalization of more comprehensive results. This creates incentives to change a proof’s structure to maximize the use of these tools, potentially resulting in less readable and “human looking” proofs.
@@
  • Non-trivial data duplication: In mathematics, it is common that one theoretical concept has several formal representations. For example, the concept of a graph can be described formally in multiple ways. Formal libraries typically feature a variety of representations for the same mathematical concept, some being almost exact copies of each other. Such near-duplication can lead to downstream issues when the library is both used as training and test data in machine learning approaches.
@@
  • Standardized interaction modes. (Analogous to natural-language, but with different characteristics) Current formal benchmarks focus on proving theorems in isolation or translating individual statements into formal code. These tasks do not capture all aspects of advanced mathematical practice.
@@
In the next section, Section 4, we discuss how certain aspects of the proof creation can be mapped to datasets. We argue that such datasets that describe intermediate stages that mathematicians go through in the long process of devising proofs are potentially highly valuable to support mathematical copilots that assist humans in these tasks. The overt focus of current datasets on result (the final proof), rather than the process that led to that proof, severely diminishes their usefulness for training LLMs to be used as mathematical copilots. While from a perspective of pure proof search, results-based datasets are justified, it is unlikely that mathematics will soon be reduced purely to proof search, as there are several mathematical activities (such as finding the “right” set of definitions, that makes a mathematical theory clear and easy to follow) that are not solely about proof search. Such datasets that describe these intermediate stages do not exist at all currently! The GHOSTS dataset, to our knowledge, is the only one that preliminarily investigates how well LLMs can engage in literature search. Thus, in Section 4.1, we take a more holistic level to describe how mathematical workflow might be described by datasets. This problem is essential, as a lot of mathematical activity is based on chaining together complex workflows. There is currently no clear way of measuring the advances of AI systems in tackling the full spectrum of mathematical workflows. Such data about intermediate stages in the proof discovery process is often not explicitly contained in textbooks and articles – the data sources on which current LLMs are trained. In Section 4.2, novel ways of capturing data that may contain this information are described. Finally, in Section 4.3, we discuss a stricter proof structure called a “motivated proof”, requiring the author to explain the origin behind each step of a proof. We then argue for an evaluation benchmark enforcing this proof structure from the perspectives of end-user utility and evidence of LLM reasoning rather than LLM memorization (coming up with a motivated proof, given an unmotivated one, requires reasoning and deeper proof understanding) and speculate on paths toward designing such a benchmark.
Lastly, in Section 5, we provide an extension of the well-known “datasheets for dataset” questionnaires (Gebru et al., 2021), specifically aimed at mathematical datasets, as these come with their own specific characteristics, not covered by the original datasheets. We hope that our questionnaire facilitates better cross-disciplinary collaboration and alignment between AI practitioners and mathematicians.
We note that in this article we do not focus on datasets that are designed primarily to be consumed by the current LLM technology stack, for example, various instruction tuning datasets, such as the OpenMathInstruct datase (Toshniwal et al., 2024). These are often based on existing datasets (GSM8K and MATH in the case of OpenMathInstruct) and thus could be described as derived datasets. Rather, our focus is on assessing how much and how well primary datasets describe various aspects of doing mathematics.
@@
Commoncore / MultiArith Roy and Roth (2015) 600 Basic arithmetic Word problems with two steps
MAWPS Koncel-Kedziorski et al. (2016) 3320 Basic arithmetic and algebra Word problems requiring systems of up to two linear equations (includes MultiArith)
Math23K Wang et al. (2017) $23161$ Basic arithmetic Word problems in Chinese
ASDiv Miao et al. (2020) $2305$ Basic arithmetic Word problems (selected from the internet)
SVAMP Patel et al. (2021) $1000$ Basic arithmetic Word problems requiring up to two steps (derived from ASDiv)
GSM8K Cobbe et al. (2021) $8500$ Basic arithmetic Word problems requiring 2-8 steps to solve
GSM8K-symbolic Mirzadeh et al. (2024) $5000$ Basic arithmetic Variants of GSM8K problems using symbolic templates
@@
Dolphin18K Huang et al. (2016) $18460$ Basic arithmetic and algebra Problems and answers from the mathematics category of Yahoo! Answers
AQuA Ling et al. (2017) $100000$ Basic arithmetic and algebra Algebraic multiple-choice word problems (similar to GMAT and GRE)
MATH Hendrycks et al. (2021) $12500$ Problem-solving Problems from high school math competitions with step-by-step solutions
NumGLUE Mishra et al. (2022a) $101835$ Arithmetic reasoning 8 different mathematical reasoning tasks (including previous datasets).
GAOKAO (Math) Zhang et al. (2023b) $844$ Various high school math topics Questions from the Chinese College Entrance Exam (single-choice, cloze, open-ended)
Hungarian High School Finals Paster (2023) $33$ Problem-solving Problems from the 2023 Hungarian National High School Finals in mathematics
MATH 401 Yuan et al. (2023) $401$ Advanced arithmetic and computation Arithmetic expressions and numerical results
TAL-SCQ5K TAL Education Group (2023) 5000 Various Multiple-choice competition questions from elementary to high school levels in English and Chinese
AGIEval (Math) Zhong et al. (2024) $1943$ Various Problems (in English and Chinese) derived from college admission tests (GRE, Gaokao, SAT) and high school competitions (from MATH and AQuA)
@@
TheoremQA Chen et al. (2023) $800$ Various (including algebra, number theory, graph theory, information theory) Questions (from the internet and textbooks) and expert answers based on $354$ theorems
ARB Sawada et al. (2023) $234$ Various Problems from university exams and competitions at the undergraduate level
@@
NaturalProofs Welleck et al. (2021) $25271$ Various (focus on algebraic geometry) Theorems with proofs and reference graphs from ProofWiki, the Stacks project, and textbooks.
GHOSTS Frieder et al. (2023a) $709$ Various Problem-solving, theorem-proving, and retrieval tasks derived from textbooks, StackExchange, and existing datasets (MATH and Symbolic integration)
FrontierMath Glazer et al. (2024) $>35$ (unknown) Various Challenging problems with automated verification created by mathematicians
@@
miniF2F Zheng et al. (2022) $488$ Problem-solving, algebra, number theory Formalized Olympiad-type problems from AIME, AMC, IMO, as well as high school and undergraduate courses (also derived from MATH)
OlympiadBench He et al. (2024) $8476$ Various Problems and solutions from mathematics and physics competitions and the Chinese College Entrance Exam (in English or Chinese and partially with images)
IMO Small Challenge Frieder et al. (2024) $100$ Combinatorics Selected problems and solutions from IMO and BWMC
IMO-AG-30 Trinh et al. (2024) $30$ Geometry Plane Euclidean geometry problems from IMO in JGEX formal language
@@
Mathematics Saxton et al. (2019) Generated Arithmetic, algebra, probability, calculus Sequential questions generated with modular structure
MMLU (Math) Hendrycks et al. (2020) $960$ Elementary, high school, and college mathematics, and abstract algebra Multiple-choice questions collected from online sources ranging from elementary to college (similar to GRE) mathematics
INT Wu et al. (2020) Generated Inequalities, General Mathematics Generated formalized (in-)equality theorems based on ordered field axioms
AMPS Hendrycks et al. (2021) $\approx 5100000$ (Linear) algebra, calculus, statistics, geometry, and number theory Problems and step-by-step solutions from Khan Academy ($>100$k) and generated using Mathematica ($\approx 5$M)
Lila Mishra et al. (2022b) 132239 Arithmetic, (linear) algebra, calculus, statistics, geometry, and number theory Derived from $20$ existing datasets (including NumGLUE, SVAMP, MultiArith, Dolphin18K, Mathematics, AMPS, GSM8K, MATH) with numbers or expressions as outputs
CMMLU (Math) Li et al. (2023) $499$ Various Multiple-choice questions (in Chinese) from elementary to college level collected from freely available resources

### v2

@@ 2 An Overview of Data-Related Issues
Several of these issues are also present for formal mathematics. But benchmarks around formal mathematics frequently suffer from distinct issues than those listed above; for instance, while datasets for formal mathematics come at a range of difficulties, from competition-level problems (zheng2021minif2f) to undergraduate textbooks (azerbayev2023proofnet), to some results that are at the mathematical state-of-the-art (scholze2022liquid; paulson2022schemes), formalized mathematics evaluation can be easily carried out in a safe and automated manner since proofs can be checked by the ITP, rendering the issue of trusted automatic evaluation irrelevant. On the other hand, the following dimensions are problematic for datasets related to formal mathematics:
@@
  • Tool misalignment: Various automation tools exist that aid in proving small, intermediate lemmas arising in the formalization of more comprehensive results. This creates incentives to change a proof’s structure to maximize the use of these tools, potentially resulting in less readable and “human looking” proofs. This misalignment often manifests as proof-script scaffolding, such as broad tactics applied repeatedly, large automation argument lists, or redundant steps that make proofs brittle and non-idiomatic. Since such redundancy is frequently detectable by linters (e.g. unused simp arguments), it can be treated as dataset metadata in corpora that include a raw view for realism along with a linter-cleaned view that better reflects canonical proof structure.
@@
  • Non-trivial data duplication: In mathematics, it is common that one theoretical concept has several formal representations. For example, the concept of a graph can be described formally in multiple ways. Formal libraries typically feature a variety of representations for the same mathematical concept, some being almost exact copies of each other. Such near-duplication can lead to downstream issues when the library is both used as training and test data in machine learning approaches. Duplication also arises at the level of lemma interfaces: when a library lacks a lemma in the form needed, developers introduce thin wrapper lemmas or local variants to patch an API gap. These near-duplicates are often engineering artifacts rather than conceptual novelty, and they can inflate the overlap between training and evaluation in ways that are invisible to string-based de-duplication.
@@
  • Standardized interaction modes: (Analogous to natural-language, but with different characteristics) Current formal benchmarks focus on proving theorems in isolation or translating individual statements into formal code. These tasks do not capture all aspects of advanced mathematical practice. In practice, formalization is dominated by interactive patterns that are mostly absent from current benchmarks, including iterative repair (compile $\rightarrow$ inspect diagnostics $\rightarrow$ patch), systematic search for lemmas or imports, and refactoring for readability and performance. For example, Lean supports interactive lemma search via exact?, apply?, and rw?, which reflect a workflow that interleaves searching and proving, rather than proof synthesis from scratch.
  • Scope of formalization and handling of incomplete proofs: Formal-math corpora vary substantially in scope and granularity. Many benchmarks emphasize short, isolated theorems or textbook-style exercises (zheng2021minif2f; azerbayev2023proofnet), whereas large-scale formalization projects capture long-form arguments that unfold across hundreds of intermediate lemmas before reaching a major theorem (scholze2022liquid; paulson2022schemes). This matters for dataset construction and evaluation: models trained or tested only on short lemmas may appear strong while still failing to manage the dependency structure, planning, and bookkeeping required for long developments.
    A second, closely related issue is incompleteness during development. In Lean, unfinished results are often recorded using sorry, and in Coq using Admitted; these placeholders accelerate library growth but may render a corpus logically incomplete (and, in the limit, inconsistent). If such items are included unmarked in training data, a model observes theorems treated as “proven” without ever seeing a proof term, which can distort learning and confound evaluation. This phenomenon is common in actively developed libraries; one example is SciLean (SciLean). Consequently, formal-math datasets should explicitly document how placeholders are handled (e.g., excluded, retained with flags, or separated into a distinct split), and should report basic completeness statistics (e.g., number or fraction of theorems containing placeholders, and where they occur in dependency graphs).
    Finally, to better reflect real mathematical practice, there is value in long-form curated corpora where an entire paper or chapter is represented as a structured development: (1) statements, (2) complete machine-checked proofs when available, (3) dependency metadata, and (optionally) (4) high-level commentary about proof strategy and lemma roles. Even when full completion is infeasible, clearly distinguishing finished proofs from partial ones enables more meaningful training objectives (e.g., proving the remaining sorry lemmas) and more reliable benchmarks.
  • Version fragility and tactic stability: Formal proofs are executable artifacts whose replay can be brittle under prover or library drift. A common failure mode is tactic instability. For instance, in Lean a “flexible” tactic such as simp changes the goal in a context-dependent way, and downstream steps silently assume a particular post-simp shape (e.g. after simplification one must start a calc chain from the actual goal state, not the expected form).
  • Hidden context dependence (imports, local instances, and attribute state): Even with a fixed statement, the behavior of a proof script can depend on surrounding environment, including imports, local instances, and attribute state (e.g. simp lemmas). In Lean, instance “pollution” can occur when local bindings introduce alternate instances (such as multiple MeasurableSpace structures), causing later lines to elaborate against an unintended instance unless the intended instance is explicitly pinned.
  • Resource sensitivity (heartbeats and elaboration/typeclass search): Correctness is not cleanly separable from performance: proofs can fail due to heartbeat limits or timeouts caused by elaboration or typeclass synthesis rather than mathematical content. This is especially salient in large developments where small refactors or additional instances can cause elaboration to take much longer, possibly more than the previous heartbeat limit.
  • Redundant scaffolding and non-canonical proof artifacts: Formal corpora often contain redundant tactic steps and linter-detectable redundancy (such as unused automation arguments, no-op steps) that are harmless for one-off development but distort the signal for training and evaluation. For instance, a tactic may already close the goal, making subsequent steps ill-typed (e.g. leading to a Lean error “no goals to be solved”). A useful dataset practice is to include both a raw corpus and a lint-cleaned canonicalized view.
  • Toolchain and automation: Many formal corpora are now partially produced by automated or semi-automated methods—ranging from autoformalization tools that attempt to translate informal text into formal proofs (jiang2022draft), to specialized rewriting tactics, to code-generation scripts that unify definitions across libraries. Current datasets may not have enough detail about which automated methods were used and how their outputs were curated. A related form of provenance is linter metadata: Mathlib-style linters flag redundant arguments, unstable patterns, and stylistic violations, and recording which lints were applied (or waived) can help downstream users understand how canonical the resulting proof scripts are. Similarly, explicit set_option declarations (e.g. heartbeat or recursion-depth overrides) should be treated as dataset metadata, since proofs that require non-default resource limits may behave differently under standard evaluation harnesses.
  • Relation of dataset to existing formal libraries: Many formal datasets build upon large existing libraries (e.g. mathlib in Lean). If a new dataset reuses or extends such a library, overlaps or version discrepancies can arise. Two projects may formalize the same theorem in slightly different ways, or rely on different versions of the library’s foundational definitions. Version fragility is particularly acute for automation-heavy proofs, as changes to typeclass instance priorities, simp-lemma sets, or default tactic behavior can cause a proof that worked in one commit to fail silently or timeout in the next, even when the underlying mathematics is unchanged.
  • Unclear benchmarking protocols: Formal mathematics datasets are frequently used as benchmarks for automated theorem proving and related model-based formal reasoning. However, papers often report a single success rate without specifying the task definition and evaluation protocol in enough detail to make results comparable.
    At minimum, a formal-math benchmark should state which task is being evaluated. Common variants include: (i) theorem proving from scratch (given a statement and imports, produce a proof that checks), (ii) tactic prediction (given a proof state, predict the next tactic or a short tactic script), (iii) proof repair or hole filling (complete a partially written proof term or script), (iv) premise selection (select a set of lemmas from a library that suffices to close a goal), and (v) autoformalization (translate an informal statement into a formal one, optionally followed by proving it).
    Even within a single task label, the protocol can change difficulty by orders of magnitude. For reproducibility, benchmark authors should specify:
    • Input specification: what the model receives (the statement only, or also local context such as imported modules, available lemmas, and current proof state), what implicit environment is assumed (active typeclass instances, scoped attributes, and simp-lemma sets), and any context limits (number of retrieved premises, maximum file size, or maximum tokens).
    • Output specification: what counts as a valid submission (a full proof script, a proof term, a sequence of tactics, or a patch to an existing file).
    • Success criterion: whether the artifact must be accepted by the checker without placeholders (e.g., no sorry/admit), and what counts as failure (type error, timeout, unsolved goals).
    • Tooling and automation: which tactics and automation are permitted (e.g., rewriting and simplification tactics, SMT calls, hammer-style tools), and whether tool use is counted as part of the model or treated as an external oracle. If multiple settings are relevant, separate tracks (for example, “with automation” and “restricted automation”) avoid conflating results.
    • Resource constraints: wall-clock time per goal, memory limits, maximum number of proof attempts, and any search limits (tactic budget, backtracking depth). Limits that are internal to the prover matter as well: heartbeat budgets, typeclass-synthesis depth, and recursion limits can all cause a correct proof script to fail silently or timeout when replayed under different settings. These choices interact strongly with inference-time compute and must be reported.
    • Environment pinning: the exact proof assistant version and library commit (for example, Lean + mathlib revision, or Isabelle + AFP snapshot), since minor version changes can make proofs succeed or fail. An evaluation harness that replays proofs in a fixed environment reduces ambiguity.
    • Metrics: beyond a single success rate, report pass@$k$ when multiple attempts are allowed, average time to solution, and optional secondary measures such as proof length or number of tactic steps, since these affect usability.
    Without these details, two papers can report the same score while solving different tasks under different constraints. Explicit protocols and pinned environments enable reliable comparison across models and make it possible to interpret gains as improvements in formal reasoning rather than changes in tooling or evaluation setup.
2.0.1 Mixed-mode datasets for tool use
Many candidate mathematical copilots are not text-in, text-out systems. They are trained and evaluated in settings where a model is expected to call external tools, such as a computer algebra system, a numerical solver, a proof assistant, a retrieval system, or a specialized symbolic model. Tool-integrated systems already appear in work on tool-integrated reasoning for math (for example, models that generate and execute Python code as part of the solution process (gou2023tora)).
A mixed-mode dataset records these interactions explicitly. Rather than storing only a problem statement and a final response, each datapoint is a short trajectory consisting of interleaved representations, for example: (i) natural-language problem context, (ii) intermediate formal objects (expressions, goals, proof states, or formal statements), (iii) tool calls with arguments, (iv) tool outputs (for example, the result of executing code, a CAS simplification, or a proof assistant error message), and (v) a final response that integrates the tool outputs into a mathematically coherent explanation. This structure supports training objectives beyond next-token prediction, including learning when to call a tool, how to translate between representations, and how to recover from tool failures (timeouts, type errors, missing imports, or non-terminating proof searches).
A special case of mixed-mode data consists of datasets pairing informal and formal mathematics. These datasets support autoformalization and naturalization, as well as theorem proving conditioned on an informal statement. Examples include miniF2F (zheng2021minif2f) and ProofNet (azerbayev2023proofnet). More recent efforts scale paired informal-formal corpora to substantially larger sizes (ying2024lean), which matters for training rather than only evaluation. In these settings, the tool is the proof assistant itself: correctness can be checked by verifying the generated formal artifact, while intermediate feedback takes the form of proof-state information and type errors, which can be recorded in the dataset.
Mixed-mode datasets introduce evaluation and reproducibility issues that do not arise in static text-only benchmarks. Tool outputs depend on the toolchain (version, library state, and sometimes randomness). A dataset intended for benchmarking should therefore specify the tool interface and pin versions (for example, by commit hashes for proof assistant libraries), and provide an evaluation harness that can replay tool calls in a controlled environment. Without this, differences in tool versions can change whether a proof checks or whether a symbolic computation succeeds, making scores hard to compare across papers.
Recommendations.
For mixed-mode datasets used for training or benchmarking, we recommend (1) representing each datapoint as an explicit tool-interaction trace, including tool inputs and outputs; (2) recording the toolchain in sufficient detail to replay the trace (versions, dependencies, and execution settings); and (3) separating tasks where tool use is optional from tasks where tool use is required, so that evaluation can distinguish mathematical reasoning from tool invocation competence. Mixed-mode datasets also interact with the contamination problem: if tool outputs encode the final answer too directly, a model can learn to pattern-match on tool outputs rather than learning to select and use tools effectively. This motivates evaluation protocols that check both the final answer and the correctness of the tool trace itself (for example, by re-executing code or re-checking formal proofs).
Issue NL Train NL Bench Formal Train Formal Bench
Difficulty calibration and coverage X X X X
Binary scoring / success-only reporting X X X X
Lack of trusted automatic evaluation X X
Standardized interaction modes (Q-A, isolated lemmas) X X X X
Contamination and leakage X X X X
Binned progress reporting (e.g., Sankey-style) X X
Tool misalignment from automation / tactic availability X X
Non-trivial duplication (multiple representations / near-copies) X X
Incomplete proofs and placeholders (sorry, Admitted) X X
Toolchain and automation provenance (scripts, autoformalization, tactics) X X
Dependence on external libraries and versioning (commits, environments) X X
Unclear benchmarking protocol (timeouts, allowed tools, environment pinning) X X
Table 2: Issue matrix for datasets used as training corpora (Train) and as benchmarks (Bench), separated by natural language (NL) and formal mathematics. “X” indicates the issue commonly or sometimes affects that use case. For formal mathematics, proof checking provides automated correctness, so the row “Lack of trusted automatic evaluation” is typically not applicable.
@@ 2.0.1 Mixed-mode datasets for tool use
In the next section, Section 4, we discuss how certain aspects of the proof creation can be mapped to datasets. We argue that such datasets that describe intermediate stages that mathematicians go through in the long process of devising proofs are potentially highly valuable to support mathematical copilots that assist humans in these tasks. The overt focus of current datasets on result (the final proof), rather than the process that led to that proof, severely diminishes their usefulness for training LLMs to be used as mathematical copilots. While from a perspective of pure proof search, results-based datasets are justified, it is unlikely that mathematics will soon be reduced purely to proof search, as there are several mathematical activities (such as finding the “right” set of definitions, that makes a mathematical theory clear and easy to follow) that are not solely about proof search. Such datasets that describe these intermediate stages do not exist at all currently! The GHOSTS dataset (frieder2023mathematical), to our knowledge, is the only one that preliminarily investigates how well LLMs can engage in literature search. Thus, in Section 4.1, we take a more holistic level to describe how mathematical workflow might be described by datasets. This problem is essential, as a lot of mathematical activity is based on chaining together complex workflows. There is currently no clear way of measuring the advances of AI systems in tackling the full spectrum of mathematical workflows. Such data about intermediate stages in the proof discovery process is often not explicitly contained in textbooks and articles – the data sources on which current LLMs are trained. In Section 4.3, novel ways of capturing data that may contain this information are described. Finally, in Section 4.4, we discuss a stricter proof structure called a “motivated proof”, requiring the author to explain the origin behind each step of a proof. We then argue for an evaluation benchmark enforcing this proof structure from the perspectives of end-user utility and evidence of LLM reasoning rather than LLM memorization (coming up with a motivated proof, given an unmotivated one, requires reasoning and deeper proof understanding) and speculate on paths toward designing such a benchmark.
We note that in this article we do not focus on datasets that are designed primarily to be consumed by the current LLM technology stack, for example, various instruction tuning datasets, such as the OpenMathInstruct dataset (toshniwal2024openmathinstruct). These are often based on existing datasets (GSM8K and MATH in the case of OpenMathInstruct) and thus could be described as derived datasets. Rather, our focus is on assessing how much and how well primary datasets describe various aspects of doing mathematics.
@@
Commoncore / MultiArith roy2015solving 600 Basic arithmetic Word problems with two steps
MAWPS koncel2016mawps 3320 Basic arithmetic and algebra Word problems requiring systems of up to two linear equations (includes MultiArith)
Math23K wang2017deep $23161$ Basic arithmetic Word problems in Chinese
ASDiv miao2020diverse $2305$ Basic arithmetic Word problems (selected from the internet)
SVAMP patel2021nlp $1000$ Basic arithmetic Word problems requiring up to two steps (derived from ASDiv)
GSM8K cobbe2021training $8500$ Basic arithmetic Word problems requiring 2-8 steps to solve
GSM8K-symbolic mirzadeh2024gsm $5000$ Basic arithmetic Variants of GSM8K problems using symbolic templates
@@
Dolphin18K huang2016well $18460$ Basic arithmetic and algebra Problems and answers from the mathematics category of Yahoo! Answers
AQuA aquarat2017ling $100000$ Basic arithmetic and algebra Algebraic multiple-choice word problems (similar to GMAT and GRE)
MATH hendrycks2021measuring $12500$ Problem-solving Problems from high school math competitions with step-by-step solutions
NumGLUE mishra2022numglue $101835$ Arithmetic reasoning 8 different mathematical reasoning tasks (including previous datasets).
GAOKAO (Math) zhang2023evaluating $844$ Various high school math topics Questions from the Chinese College Entrance Exam (single-choice, cloze, open-ended)
Hungarian High School Finals testing_language_models_on_a_held_out_high_school_national_finals_exam $33$ Problem-solving Problems from the 2023 Hungarian National High School Finals in mathematics
MATH 401 yuan2023well $401$ Advanced arithmetic and computation Arithmetic expressions and numerical results
TAL-SCQ5K tal-scq5k 5000 Various Multiple-choice competition questions from elementary to high school levels in English and Chinese
AGIEval (Math) zhong2024agieval $1943$ Various Problems (in English and Chinese) derived from college admission tests (GRE, Gaokao, SAT) and high school competitions (from MATH and AQuA)
@@
TheoremQA chen2023theoremqa $800$ Various (including algebra, number theory, graph theory, information theory) Questions (from the internet and textbooks) and expert answers based on $354$ theorems
ARB sawada2023arb $234$ Various Problems from university exams and competitions at the undergraduate level
@@
NaturalProofs welleck2021naturalproofs $25271$ Various (focus on algebraic geometry) Theorems with proofs and reference graphs from ProofWiki, the Stacks project, and textbooks.
GHOSTS frieder2023mathematical $709$ Various Problem-solving, theorem-proving, and retrieval tasks derived from textbooks, StackExchange, and existing datasets (MATH and Symbolic integration)
FrontierMath glazer2024frontiermath >35>35 (unknown) Various Challenging problems with automated verification created by mathematicians
@@
miniF2F zheng2021minif2f $488$ Problem-solving, algebra, number theory Formalized Olympiad-type problems from AIME, AMC, IMO, as well as high school and undergraduate courses (also derived from MATH)
OlympiadBench he2024olympiadbench $8476$ Various Problems and solutions from mathematics and physics competitions and the Chinese College Entrance Exam (in English or Chinese and partially with images)
IMO Small Challenge frieder2024imo $100$ Combinatorics Selected problems and solutions from IMO and BWMC
IMO-AG-30 trinh2024solving $30$ Geometry Plane Euclidean geometry problems from IMO in JGEX formal language
@@
Mathematics saxton2019analysing Generated Arithmetic, algebra, probability, calculus Sequential questions generated with modular structure
MMLU (Math) hendrycks2020measuring $960$ Elementary, high school, and college mathematics, and abstract algebra Multiple-choice questions collected from online sources ranging from elementary to college (similar to GRE) mathematics
INT wu2020int Generated Inequalities, General Mathematics Generated formalized (in-)equality theorems based on ordered field axioms
AMPS hendrycks2021measuring $\approx 5100000$ (Linear) algebra, calculus, statistics, geometry, and number theory Problems and step-by-step solutions from Khan Academy (>100>100k) and generated using Mathematica ($\approx 5$M)
Lila mishra-etal-2022-lila 132239 Arithmetic, (linear) algebra, calculus, statistics, geometry, and number theory Derived from $20$ existing datasets (including NumGLUE, SVAMP, MultiArith, Dolphin18K, Mathematics, AMPS, GSM8K, MATH) with numbers or expressions as outputs
CMMLU (Math) li2023cmmlu $499$ Various Multiple-choice questions (in Chinese) from elementary to college level collected from freely available resources